# Less is KEN: a Universal and Simple Non-Parametric Pruning Algorithm for Large Language Models

### v1

@@ 6.2 Compression values
Model Total params Original file size # trainable params Compressed file size (Model + support dict) BLOOM 1.72B 7,055 MB 664M 3,071 MB (2,923 + 148) 442M 2137 MB (2,013 + 124) BLOOM 560M 2,294 MB 429M 2,084 MB (1,956 + 128) 386M 1,842 MB (1,731 + 111) BERT 109M 438 MB 80M 358 MB (320 + 38) 57M 260.2 MB (228 + 32.2) DistilBERT 66M 266 MB 51M 231.4 MB (203 + 28.4) 36M 165 MB (145 + 20) DeBERTa 138M 555 MB 107M 476.3 MB (428 + 48.3) 76M 348.4 MB (306 + 42.4) Ernie 109M 438 MB 80M 356.9 MB (320 + 36.9) 57M 260.3 MB (228 + 32.3) Electra 33M 134 MB 14M 67.01 MB (59.1 + 7.91) 9M 42.58 MB (35.5 + 7.08)
One of the primary objectives of KEN is to significantly reduce the overall size of transformer models, including their file sizes. To accomplish this goal, KEN leverages a subnetwork comprising only -trained parameters, allowing it to be saved and then injected into its pre-trained counterpart. This process requires a support file, like a dictionary, that specifies the precise location of each saved parameter within the pre-trained model. To ensure a fair comparison between the original and compressed model sizes, the compressed model is saved using the same techniques and format as the original model, guaranteeing consistent results. For each model, two compressed versions are generated, employing both high and low values.
As shown in Tab. 5, both versions of the compressed models exhibit substantial memory savings, with the size of the compressed model directly proportional to the number of saved parameters. Specifically, models saved using a high value, and thus closely mirroring the structure of the unpruned model, conserve 100 MB per model. This value further increases as the number of trained parameters saved diminishes. The support dictionary for parameter injection, stored using the Lempel-Ziv-Markov chain data compression algorithm, has an insignificant impact on the model final weight, which remains significantly smaller than the original. Furthermore, the time required to load the injected parameters into the pre-trained model is linear with the transformer architecture and the compression employed.
6.3 KEN
KEN is a visualization tool that provides a clear understanding of the composition of matrices after the application of KEN pruning step. It offers various views to explore the pruned model, including:
@@ 6.3 KEN
Single Matrix View: It displays only the retained parameters, leaving the pruned ones blank (Fig. 2).
@@
Neighbor Count View: It visualizes the number of non-zero neighbors (horizontally and vertically) for each point in a given matrix.
@@
Layer-wise View: This iterative view applies the previous two views to each matrix per model layer.
@@ 7 Conclusions
In this paper, we presented KEN, a novel non-architecture-specific pruning algorithm that leverages KDE to construct an abstraction of the parameter distribution and selectively retain a finite subset of parameters while resetting the rest to their pre-trained values. Our extensive evaluations on seven diverse transformer models demonstrate that KEN consistently achieves remarkable compression rates, reducing unnecessary parameters by a minimum of 25% up to 70% on some models, without compromising model performance. Moreover, by leveraging the KEN core idea, is possible to store only the subnetwork of -trained parameters, leading to significant memory savings. We also present KEN: the KEN visualizer that provides insights into the algorithm operation. KEN reveals that KEN uniformly selects parameters across matrices, hindering cluster formation. With KEN we demonstrate how a simple, non-parametric strategy commonly used in statistics can be adopted for model pruning to obtain excellent results in terms of compression and performance.
@@ 8 Limitations
One of the major limitation of KEN is its computational efficiency, particularly when analyzing large models. Although KEN excels at producing detailed distributions using large values, this comes at the cost of increased processing time. The computational effort increases linearly with the size of the model matrix, the number of model layers, and the chosen value. It is important to note that this performance impact mainly affects the parameter selection phase and does not significantly affect the saving or loading of compressed models.
Furthermore, although our paper focused on the sequence classification task to ensure complete and comparable results, preliminary unpublished experiments demonstrate the effectiveness of KEN in other tasks. Future work will explore its broader applicability and address potential optimizations for large-scale scenarios.
@@ Appendix A How to prove the importance of selected parameters
To assess the effectiveness of KEN core idea, which involves selecting parameters based on their distribution using Kernel Density Estimation (KDE), we conducted parallel experiments. In these experiments, KEN randomly chose parameters to either retain or reset to their pre-trained values. This randomized approach allowed us to compare KEN KDE-based selection strategy against random parameter pruning.
@@
The results, illustrated in Fig. 5, consistently show that KEN outperforms its random counterpart. KEN achieves a lower error rate and a smaller performance gap at reasonable compression levels. It is important to note that, in all cases and for all models examined, there exists a threshold value beyond which the model performance inevitably declines. The KEN algorithm effectively compresses models while preserving high performance and minimizing error rates. However, if the reset parameters exceed a certain threshold (specific to each model), its performance suffers a catastrophic decline. When random values are used, this threshold is reached earlier, resulting in a larger performance gap and higher error rate. Nevertheless, the upper limit obtained with random selection is always lower than or equal to the average value obtained with KEN.
@@
Dataset BLOOM BLOOM Bert DistilBert DeBERTa Ernie Electra trec 61.46% 23.26% 26.55% 23.45% 22.84% 26.55% 55.94% rotten_tomatoes 69.17% 24.80% 26.55% 34.39% 44.88% 42.29% 55.94% hate_speech_offensive 61.46% 23.26% 26.55% 34.39% 22.84% 26.55% 55.94% hate_speech18 61.46% 23.26% 26.55% 23.45% 33.86% 31.80% 64.75% scicite 61.46% 23.26% 37.05% 28.92% 22.84% 31.80% 55.94% ade_corpus_v2 69.17% 24.80% 52.78% 45.32% 44.88% 63.28% 73.56% amazon_reviews_multi 69.17% 24.80% 31.80% 34.39% 22.84% 31.80% 55.94% poem_sentiment 74.31% 26.34% 58.03% 45.32% 22.84% 47.54% 73.56% tweet_eval-emoji 74.31% 23.26% 63.28% 23.45% 44.88% 79.02% 55.94% tweet_eval-hate 61.46% 23.26% 26.55% 61.73% 44.88% 47.54% 55.94% tweet_eval-irony 61.46% 23.26% 26.55% 23.45% 22.84% 26.55% 64.75% tweet_eval-offensive 61.46% 23.26% 26.55% 34.39% 28.35% 31.80% 55.94% tweet_eval-femminist 61.46% 23.26% 26.55% 39.05% 22.84% 37.05% 64.75%
@@ Appendix B Additional results
In this appendix, we show additional results obtained using KEN not shown in Tab. 2. Tab. 7 provides a comprehensive overview of all datasets analyzed in the paper, while Tab. 6 displays the additional results included. Unlike Tab. 2, Tab. 6 focuses on the highest percentage of reset parameters for each model on each dataset where KEN F1-weighted score matches or surpasses the performance of the original unpruned model. This highlights the exceptional compression capabilities of KEN, enabling it to achieve comparable or even improved performance while significantly reducing the model parameter count.
@@
Appendix C KEN examples
The goal of KEN is to generate visual representations of the pruning results obtained from KEN. In this example, we highlight the key matrices of layers 0 and 12 of a BERT model, trained on the glue-sst2 dataset. The visualizations reveal the parameters selected by KEN and their respective neighbor counts, as discussed in Sec. 6.3.
In this experiment, we used BERT instead of other models analyzed in the paper, because it performed remarkably well at both low and high values during the testing phase (shown in Tab. 2 and Tab. 6). To fully examine the evolution of parameter selection patterns, we used three different values, representing different degrees of selected parameters. This allowed us to observe how these parameters changed as the amount of parameter resetting increased.
From Fig. 6 and Fig. 7, it is clear that in all configurations and in all analyzed layers, the distribution of points in each row of each matrix is relatively uniform and does not deviate into a distinct, disconnected cluster. Additionally, the number of non-zero neighbors for each point is quite uniform even as the value varies.

### v2

@@ 6.2 Compression values
One of KEN main goals, is to significantly reduce the overall size of transformer models, including their file sizes. To achieve this, KEN leverages a subnetwork comprising only -trained parameters, allowing it to be saved and then injected into its pre-trained counterpart. This process requires a support file, like a dictionary, that specifies the precise location of each saved parameter within the pre-trained model. To guarantee an unbiased comparison between the original and compressed model sizes, the compressed one is saved using identical techniques and format as the original model. For each transformer analyzed, two compressed versions are generated using both high and low values.
Model Total params Original file size # trainable params Compressed file size (Model + support dict) BLOOM1B7 1.72B 7,055 MB 664M 3,071 MB (2,923 + 148) 442M 2137 MB (2,013 + 124) BLOOM560k 560M 2,294 MB 429M 2,084 MB (1,956 + 128) 386M 1,842 MB (1,731 + 111) BERT 109M 438 MB 80M 358 MB (320 + 38) 57M 260.2 MB (228 + 32.2) DistilBERT 66M 266 MB 51M 231.4 MB (203 + 28.4) 36M 165 MB (145 + 20) DeBERTa 138M 555 MB 107M 476.3 MB (428 + 48.3) 76M 348.4 MB (306 + 42.4) Ernie 109M 438 MB 80M 356.9 MB (320 + 36.9) 57M 260.3 MB (228 + 32.3) Electra 33M 134 MB 14M 67.01 MB (59.1 + 7.91) 9M 42.58 MB (35.5 + 7.08)
As shown in Tab. 5, both versions of the compressed models demonstrate substantial memory savings, directly proportional to the number of saved parameters. Specifically, transformers saved using a high value, thus closely mirroring the structure of the unpruned model, conserving 100 MB per each. This value increases further as the number of trained parameters saved diminishes. The support dictionary for parameter injection, stored using the Lempel-Ziv-Markov chain data compression algorithm, has a negligible impact on the final model weight, which remains significantly smaller than the original. Additionally, the time required to load the injected parameters into the pre-trained model scales linearly with the transformer architecture and the compression employed.
6.3 KENviz
KENviz is a visualization tool that provides a clear understanding of matrices composition after the application of KEN pruning step. It offers various views to explore the pruned model, including:
@@ 6.3 KENviz
Neighbor Count View: It visualizes the number of non-zero neighbors (horizontally and vertically) for each point in a given matrix. This provides additional information about the remaining parameters and potential parameter clusters that might emerge (Fig. 5(b)).
@@
Layer-wise View: This view shows the impact of KEN on the entire model architecture. It iteratively applies the Single matrix view to the same matrix on all layers it appears. This allows a layer-by-layer comparison, revealing how the pruning of the same matrix changes in different parts of the network.
@@ 7 Conclusions
This paper introduces KEN: a novel non-architecture-specific pruning algorithm that leverages KDE to construct an abstraction of the parameter distribution and selectively retain a finite subset of them while resetting the others to their pre-trained states. Our extensive evaluations on seven different transformer models demonstrate that KEN consistently achieves remarkable compression rates, reducing unnecessary parameters by a minimum of 25% up to 70% on some models, without compromising model performance. Moreover, by leveraging the KEN core idea, is possible to store only the active subnetwork, leading to substantial memory savings.
We also present KENviz: the KEN visualizer that provides insights into the algorithm operation. KENviz reveals that KEN uniformly selects parameters across matrices, preventing parameter clusters. With KEN we demonstrate how a simple, non-parametric algorithm commonly used in statistics, can be effectively adopted for model pruning achieving excellent results in both compression and performance.
@@ 8 Limitations
One of the major limitations of KEN is its computational efficiency, especially when analyzing large models. Although KEN excels at generating rich distributions with high values, it faces a trade-off in computational efficiency, particularly for large models. Processing time scales linearly with the model size, number of layers, and chosen . This primarily impacts the parameter selection stage, not compressed model storage or loading.
Our focus on the sequence classification task ensured consistent results, but unpublished experiments suggest KEN effectiveness extends to other tasks. Future work will explore this broader applicability and investigate potential optimizations for large-scale deployments.
@@ Appendix A How to prove the importance of selected parameters
To assess the effectiveness of KEN core idea, which involves selecting parameters based on their distribution using Kernel Density Estimation (KDE), we conducted parallel experiments. In these experiments, we compared KEN against random parameter pruning. Random parameters were either retained or reset to pre-trained values. This enabled us to determine if the selection method used in KEN resulted in a better subnetwork compared to random selection
@@
The results, illustrated in Fig. 6, consistently show that KEN outperforms its random counterpart. KEN achieves a lower error rate and maintains higher performance at reasonable compression levels. However, for all models tested, a compression threshold exists beyond which performance inevitably declines. The KEN algorithm effectively compresses models while minimizing this decline, preserving high performance. Conversely, random parameter selection reaches this threshold earlier, resulting in a larger performance drop and higher error rate. Notably, the best achievable performance with random selection is always lower than or equal to the average achieved with KEN.
@@
Dataset BLOOM1B7 BLOOM560k Bert DistilBert DeBERTa Ernie Electra trec 61.46% 23.26% 26.55% 23.45% 22.84% 26.55% 55.94% rotten_tomatoes 69.17% 24.80% 26.55% 34.39% 44.88% 42.29% 55.94% hate_speech_offensive 61.46% 23.26% 26.55% 34.39% 22.84% 26.55% 55.94% hate_speech18 61.46% 23.26% 26.55% 23.45% 33.86% 31.80% 64.75% scicite 61.46% 23.26% 37.05% 28.92% 22.84% 31.80% 55.94%† ade_corpus_v2 69.17% 24.80% 52.78% 45.32% 44.88% 63.28% 73.56% amazon_reviews_multi 69.17% 24.80% 31.80% 34.39% 22.84% 31.80% 55.94%† poem_sentiment 74.31% 26.34% 58.03% 45.32% 22.84% 47.54% 73.56% tweet_eval-emoji 74.31% 23.26% 63.28% 23.45% 44.88% 79.02% 55.94% tweet_eval-hate 61.46% 23.26% 26.55% 61.73% 44.88% 47.54% 55.94% tweet_eval-irony 61.46% 23.26% 26.55% 23.45% 22.84% 26.55% 64.75% tweet_eval-offensive 61.46% 23.26% 26.55%† 34.39% 28.35% 31.80% 55.94% tweet_eval-femminist 61.46% 23.26% 26.55% 39.05% 22.84% 37.05% 64.75%
@@ Appendix B Additional results
This appendix presents additional results obtained using KEN that are not included in Tab. 2.
Tab. 7 provides a comprehensive overview of all datasets analyzed in the paper. In contrast to Tab. 2, Tab. 6 focuses on the specific results included in this appendix. Here, we highlight cases where KEN achieves F1-weighted scores that match or surpass the original unpruned model performance, focusing on the highest percentage of reset parameters for each model on each dataset. This emphasizes KEN remarkable compression capabilities, allowing it to achieve comparable or even improved performance while significantly reducing the model parameter count.
@@
Appendix C KENviz examples
KENviz generates visual representations of the model pruning after the KEN application. Here, we focus on key matrices from layers 0 and 12 of a BERT model trained on the glue-sst2 dataset (details in Sec. 6.1). For each layer, we present both a single matrix view and a neighbor count view, as described in Sec. 6.3.
BERT was chosen for this experiment due to its exceptional performance across a range of k values during testing (Tab. 2 and Tab. 6). To comprehensively explore how parameter selection patterns evolve, we employed three different values, representing varying degrees of parameter selection. This allowed us to observe how parameter choices shift as the amount of parameter resetting increases.
Fig. 7 and Fig. 8 consistently reveal a uniform distribution of parameters within each matrix row across all configurations and layers. This implies an absence of well-defined clusters of selected parameters. Furthermore, the number of neighbors for each parameter remains consistent regardless of the chosen value.